OpenStudy (anonymous):

Is there a quick way to simplify this? f"(x)=((x²+3)²(-36)-(-36x)(2)(x²+3)(2x))/(x²+3)⁴

OpenStudy (kainui):

Get good at algebra or use the product rule instead of the quotient rule always.

OpenStudy (zehanz):

Written as a proper formula:\[f"(x)=\frac{ (x^2+3)^2 \cdot -36--36x \cdot 2(x^2+3) \cdot 2x }{ (x^2+3)^4 }\]You see that x^2+3 is a factor of the numerator and the denominator, so it cancels:\[f"(x)=\frac{ (x^2+3) \cdot -36+36x \cdot 2 \cdot 2x }{ (x^2+3)^3 }\]Also, 36 can be factored out in the numerator :\[f"(x)=\frac{ 36(4x^2-(x^2+3))}{ (x^2+3)^3 }\]Now just simplify the numerator a little more:\[f"(x)=\frac{ 36(3x^2-3)}{ (x^2+3)^3 }=\frac{ 108(x^2-1)}{ (x^2+3)^3 }=\frac{ 108(x+1)(x-1)}{ (x^2+3)^3 }\]That's about the quickest, I'd say...

OpenStudy (anonymous):

Awesome, thats exactly what I was looking for I appreciate the help.

OpenStudy (zehanz):

YW!